Kolby’s Korndogs is looking at a new sausage system with an installed cost of $655,000. This cost will be depreciated straight-line to zero over the project’s five-year life, at the end of which the sausage system can be scrapped for $85,000. The sausage system will save the firm $183,000 per year in pretax operating costs, and the system requires an initial investment in net working capital of $35,000. If the tax rate is 22 percent and the discount rate is 8 percent, what is the NPV of this project? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Annual depreciation = 655,000 / 5 = 131,000

initial investment = Cost + NWC

initial investment = 655,000 + 35,000 = $690,000

Operating cash flow form year 1 to year 5 = (saving - depreciation)(1 - tax) + depreciation

Operating cash flow form year 1 to year 5 = (183,000 - 131,000)(1 - 22) + 131,000

Operating cash flow form year 1 to year 5 = 40,560 + 131,000

Operating cash flow form year 1 to year 5 = 171,560

Year 5 non operating cash flow = 85,000 + 35,000 - 0.22(85,000 - 0)

Year 5 non operating cash flow = 85,000 + 35,000 - 18,700

Year 5 non operating cash flow = 101,300

NPV = Present value of cash inflows - present value of cash outflows

NPV = -690,000 + 171,560 / (1 + 0.08)1 + 171,560 / (1 + 0.08)2 + 171,560 / (1 + 0.08)3 + 171,560 / (1 + 0.08)4 + 171,560 / (1 + 0.08)5 + 101,300 / (1 + 0.08)5

NPV = $63,932.41